QUESTION 4 Compute d O-8.454 3 0 -3.716 O -3.195 O -0.262 O0.979 O 3.082 3.210 O 5.703 O 6.114
Answer #1

4.6)

First Value Second Value d
2.45 0.5 -1.95
1.43 -5.34 -6.77
3.68 8.4 4.72
-3.29 4.18 7.47
-5.14 2.03 7.17
-1.49 7.44 8.93
4.44 8.1 3.66
-0.68 5.58 6.26
4.13 3.53 -0.6

dbar = 3.21

s(dbar) = 5.2330

SE = s(dbar)/sqrt(n)

= 5.2330/sqrt(9)

= 1.7443

Test Statisitcs, t = dbar/SE

= 3.21/1.7443

= 1.8402

degrees of freedom = n-1 = 8

p-value = 0.0515

As p-value is greater than the significance level of 0.05, we fail to reject the null hypothesis.